OpenStudy (anonymous):

Solve teh following equation for "x" 2sqr(x) = x-3

OpenStudy (anonymous):

square both sides get \[4x=(x-3)^2\] expand the right hand side to get \[4x=x^2-6x+9\] then put \[x^2-10x+9=0\] and solve the resulting quadratic equation by factoring

OpenStudy (anonymous):

then check your answers in the original equation and one will not work, because \[2\sqrt{x}\] is positive, not negative